# Flat quasi-Frobenius Lie superalgebras

### v1

@@ 7.3. Five-dimensional flat quasi-Frobenius Lie superalgebras
Let $(\mathfrak{b},\omega_{\mathfrak{b}})$ be the three-dimensional orthosymplectic quasi-Frobenius abelian Lie superalgebra.
@@
Since $b_{0}$ is even, we can write $b_{0}=\alpha e_{1}+\beta e_{2}$, $\alpha,\beta\in\mathbb{R}$.
@@
$\xi=\begin{pmatrix}0&0&a\\ 0&0&0\\ \alpha&\beta&0\end{pmatrix},\qquad a\neq 0,\quad\alpha,\beta\in\mathbb{R}.$
@@
Let $(\mathfrak{g},\omega)$ be a flat orthosymplectic quasi-Frobenius non-abelian Lie superalgebra of total dimension $5$. According to Theorem 6.7.  Theorem, such an algebra is an odd or even flat double extension of the three-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. By Prop. 7.3.1.  Proposition, we have:
@@
    If $\xi=0$ and $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}+f_{1}^{*}\wedge f^{*}_{1}$, where $\alpha,\beta\in\mathbb{R}$, then according to (21), there exists a basis $\{e,d,e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
@@
    $(x_{1},x_{2},x_{3},x_{4}\mid y_{1})=(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1},f_{1}).$
    Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2},x_{3},x_{4}\mid y_{5})=\left(3ad+\alpha e_{2},d+\frac{\alpha+1}{3a}e_{2},\alpha e-3ae_{1},-\frac{\alpha+1}{3a}e+e_{1},f_{1}\right).$
@@
    $x_{1},x_{2}]=x_{3}\mbox{ and }\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*}+y_{1}^{*}\wedge y_{1}^{*},$
    which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
@@
$(x_{1},x_{2}\mid y_{1},y_{2},y_{3})=(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e,f_{1}).$
Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$. ∎

### v2

@@ 7.2. Five-dimensional flat quasi-Frobenius Lie superalgebras
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be the three-dimensional orthosymplectic quasi-Frobenius abelian Lie superalgebra.
@@
Since $b_{0}$ is even, we can write $b_{0}=\alpha e_{1}+\beta e_{2}$, $\alpha,\beta\in\mathbb{K}$.
@@
$\xi=\begin{pmatrix}0&0&a\\ 0&0&0\\ \alpha&\beta&0\end{pmatrix},\qquad a\neq 0,\quad\alpha,\beta\in\mathbb{K}.$
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius non-abelian Lie superalgebra of total dimension $5$. According to Theorem 6.7.  Theorem, such a superalgebra is an odd or even flat double extension of the three-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. By Prop. 7.2.1.  Proposition, we have:
@@
    If $\xi=0$ and $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}+f_{1}^{*}\wedge f^{*}_{1}$, where $\alpha,\beta\in\mathbb{K}$, then according to (21), there exists a basis $\{e,d,e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
@@
    $(x_{1},x_{2},x_{3},x_{4},y_{1})=\left(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1},f_{1}\right).$
    Thus $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2},x_{3},x_{4},y_{5})=\left(3ad+\alpha e_{2},d+\frac{\alpha+1}{3a}e_{2},\alpha e-3ae_{1},-\frac{\alpha+1}{3a}e+e_{1},f_{1}\right).$
@@
    $x_{1},x_{2}]=x_{3}\quad\text{ and }\quad\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*}+y_{1}^{*}\wedge y_{1}^{*},$
    which implies that $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
@@
$(x_{1},x_{2}\mid y_{1},y_{2},y_{3})=\left(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e,f_{1}\right).$
Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$. ∎
Acknowledgments: We sincerely thank the referee for the valuable comments, which have improved the clarity and readability of the manuscript.
Declarations:
FUNDING STATEMENT: SB’s research was supported by GRANT: AD-065.
COMPETING INTEREST STATEMENT: the authors declare no competing interest.